OpenStudy (anonymous):

Solve the equation. 12 + 0.35x = 20.05 A. 91.5 B. 57.3 C. 2.8175 D. 23

OpenStudy (cj49):

20.05-12=?

OpenStudy (anonymous):

8.05? @cj49

OpenStudy (cj49):

x=8.05/0.35=?

OpenStudy (anonymous):

23?

OpenStudy (cj49):

thats the ans mate..

OpenStudy (anonymous):

Cool! Thanks!